A charge, q1, of +17 micro-coulombs is located at x = 0, y = 9.2 cm, a charge q2 of -42 micro-coulombs is located at x = 0, y = -3.9 cm, an unknown charge of q3, is located at x =- 3.9 cm, y = 0 cm, and a charge, q4, -6 micro-coulombs is located at x = +10.7 cm, y = 0 cm. The direction of the total electric force on q4 is 135.4 degrees from the +x axis, counter-clockwise. What is the unknown charge, q3, in micro-coulombs? If negative, include a negative sign in your answer.
Answer #1

1 07 Bo. me-t D.039m G4=-6x10-6 C Yo.7

FA = force by charge at A = k q1 q4/AD2 = (9 x 109) (17 x 10-6) (6 x 10-6)/(0.141)2 = 46.2 N

FAy = Y-component of force FA = FA Sinheta = 46.2 Sin40.7 = 30.13 N

FAx = X-component of force FA = - FA Cosheta = - 46.2 Cos40.7 = - 35.03 N

FB = force by charge at B = k q2 q4/BD2 = (9 x 109) (42 x 10-6) (6 x 10-6)/(0.114)2 = 174.5 N

FBy = Y-component of force FB = FB Sinheta' = 174.5 Sin20 = 59.7 N

FBx = X-component of force FB = FB Cosheta' = 174.5 Cos20 = 164 N

Fc = force by charge at C = k q3 q4/CD2 = (9 x 109) q3 (6 x 10-6)/(0.146)2

FCy = Y-component of force FC = 0

FCx = X-component of force FC = (9 x 109) q3 (6 x 10-6)/(0.146)2

Fy = Y-component of the net force = Fnet Sin135.4

Along the Y-direction , force equation is given as

FAy + FBy = Fy

30.13 + 59.7 = Fnet Sin135.4

Fnet = 128 N

Along the X-direction , force equation is given as

FCx + FBx + FAx = Fx

FCx + (164) + ( - 35.03) = (128) Cos135.4

FCx = - 220.11 N

magnitude of force Fc = 220.11 N

(9 x 109) q3 (6 x 10-6)/(0.146)2 = 220.11

q3 = 86.88 x 10-6 C